Three cards are drawn randomly from a deck of cards one after the other with replacement each time. Find the probability that the three cards are a 5, a heart and a king.

a. 0.0156
b. 0.125
c. 0.0015
a. 0.0192
1
Expert's answer
2020-12-24T16:18:00-0500
"P=P(5)P(Heart)P(King)=\dfrac{4}{52}(\dfrac{13}{52})(\dfrac{4}{52})"

"=\dfrac{1}{13}(\dfrac{1}{4})(\dfrac{1}{13})=\dfrac{1}{676}"

If the order is important


"P=\dfrac{1}{676}\approx0.0015"

If the order is not important, there are 6 possible variants

"3!=6""(5, H, K), (5, K, H), (H, 5, K), (H,K,5),"

"(K,5,H),(K,H,5)"

"P=6(\dfrac{1}{676})=\dfrac{3}{338}\approx0.0089"

c. 0.0015